Find the original price of a pair of shoes if the sale price is $105 after a 30 percent discount
Answer by Susan-math(40)

Let P=original price.
Then sale price = p - (.30)p=.70p
and .70p=105
so p=105/.70=150
Check: if Price was 150 and reduced by (.30)(150)=$45,
then sale price = 150-45=$105.